The sum of three consecutive odd numbers is 51. Find the numbers.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let the three consecutive odd numbers are x, x+2 and x+4.

As per the given question, the sum all the three numbers is 51

Therefore,

x+x+2+x+4=51

3x+6=51

x=(51-6)/3=45/3 =15

Hence the numbers are:

x=15

x+2=17

x+4=19

Thus, the required three consecutive odd numbers are 15,17 and 19.
Thanks